OpenStudy (anonymous):

(x^2 x x^4)^-3= first express the answer in exponential form. then evaluate the expression.

OpenStudy (anonymous):

x^-a=1/x^a

OpenStudy (gw2011):

[(x^2)(x^4)]^-3=(x^6)^-3=x^-18=1/x^18